Consider the following sample data for the relationship between advertising budget and sales for Product A:

Observation 1 2 3 4 5 6 7 8 9 10
Advertising ($) 90,000 100,000 100,000 110,000 120,000 120,000 130,000 140,000 140,000 150,000
Sales ($) 535,000 626,000 625,000 674,000 712,000 725,000 809,000 832,000 845,000 919,000

What is the slope of the "least-squares" best-fit regression line?

Please round your answer to the nearest hundredth.

Note that the correct answer will be evaluated based on the full-precision result you would obtain using Excel.
Answer #1

Solution:

x y x^2 xy
90000 535 8100000000 48150000
100 626 10000 62600
100 625 10000 62500
110 674 12100 74140
120 712 14400 85440
120 725 14400 87000
130 809 16900 105170
140 832 19600 116480
140 845 19600 118300
150000 919000 2.25E+10 1.38E+11
SUM 240960 925383 3.06E+10 1.38E+11

nn = 10,

Σx =240960

Σy =925383

  x *y = 1.38 * 1011

2 Σ = 3.06 * 1010

     1596212645452_blob.png

let a linear model

y=potßix                                                                                                                                             (1)

Σy = βο Σ1 +β, Σx                     

925383 - 10B. +24096031                                                                                                                         (2)                                                                                                                   

Multiplying by x in eqn. (1) and taking sum both side

Σxy = βοΣ +β, Σx?

1.38 * 1011 = 240960B. +3.06 * 1010B1                                                                                                                       (3)

Solving eqn. (2) , (3)

Bo = -1.9907 * 104 = -19907

B. = 4.6666

linear model is y == -19907 + 4.6666x

So slope is β1=4.6666B. = 4.6666